OpenStudy (anonymous):

Find all solutions on the interval (0,2pi) tan x sin2 x= tan x

OpenStudy (anonymous):

If tan(x)=0 then the equation works. You can also go about it this way: tan(x)sin(2x)=tan(x) Divide both sides by tan(x) sin(2x)=1 Let u=2x

OpenStudy (anonymous):

tan x(sin2x -1) =0 (DO NOT CROSS OUT THE TAN X!!!) => tan x = 0 or sin 2x = 1